A random sample of 23 components is drawn from a population size of 100. The mean

measurement of the random sample is 67.45mm and has a standard deviation of 2.93 mm.

Determine a 90% confidence interval for an estimate of the mean of the population.


1
Expert's answer
2022-02-10T16:10:27-0500
"s=\dfrac{a}{b}\sqrt{\dfrac{a}{b}}"

The critical value for "\alpha = 0.1" and "df = n-1 = 22"  degrees of freedom is "t _c = z_{1-\alpha/2; n-1} = 1.717144." The corresponding confidence interval is computed as shown below:


"CI=(\bar{X}-t_c\times\dfrac{s}{\sqrt{n}}\sqrt{\dfrac{N-n}{N-1}},"

"\bar{X}+t_c\times\dfrac{s}{\sqrt{n}}\sqrt{\dfrac{N-n}{N-1}})"

"=(67.45-1.717144\times\dfrac{2.93}{\sqrt{23}}\sqrt{\dfrac{100-23}{100-1}},"

"67.45+1.717144\times\dfrac{2.93}{\sqrt{23}}\sqrt{\dfrac{100-23}{100-1}})"

"=(66.5248, 68.3752)"

Therefore, based on the data provided, the 90% confidence interval for the population mean is "66.5248 < \mu < 68.3752," which indicates that we are 90% confident that the true population mean "\mu" is contained by the interval "(66.5248,68.3752)."